OpenStudy (anonymous):

Determine whether the system of linear equations has one and only one solution, infinitely many solutions, or no solution. 2x− y= 5 3x+ y=−6

OpenStudy (anonymous):

Add the two equations and solve for x.

OpenStudy (anonymous):

add them? im not supposed to solve?

OpenStudy (anonymous):

i meant the first one

OpenStudy (anonymous):

Adding the two equations will give you an equation in just x. Then, you can solve for x. Using that value of x, you can solve for y.

OpenStudy (anonymous):

5x+y=-1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

5x = -1 Remember, y and -y zero out.

OpenStudy (anonymous):

so now what do i do?

Directrix (directrix):

only one solution

OpenStudy (mertsj):

The two equations do not have the same slope. So they are not parallel. The two equations are not the same equation. so they are not the same line. Therefore, they intersect in 1 point and there is 1 solution.

OpenStudy (anonymous):

how do i go about finding the points they intersect at?

OpenStudy (mertsj):

Solve the system using either elimination or substitution.

OpenStudy (anonymous):

5x = -1 So, x = -1/5. Plug this value of x in one of the equations. You will get an equation in y. Solve for y.

OpenStudy (anonymous):

the fraction messes me up

OpenStudy (anonymous):

does that equal out to .5?

OpenStudy (anonymous):

is y 3?

OpenStudy (mertsj):

No. 1/5 = .2